If you know that two legs of a right triangle are 6 and 13 units long, what is the triangle's area? If necessary, round your answer to two decimal places.
Answer by macston(5194)

In a right triangle, the legs are base and height, and area =1/2bh
so:
1/2(6 units)(13 units)=area
1/2(78 sq units)=area
39 sq units=area